OpenStudy (anonymous):

Integrate (2x-1)/(3) I'm not so sure how to do this so any help would be appreciated. Thanks

zepdrix (zepdrix):

\[\Large \int\limits \frac{2x-1}{3}dx\]So that's our integral? :)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I'm not so sure how to rearrange it into a simpler form for integrating

OpenStudy (anonymous):

Using the definition of the integral?

OpenStudy (anonymous):

Or using integral properties?

OpenStudy (anonymous):

Well this is part of a question to find the area of a shaded segment

OpenStudy (anonymous):

Well do you know how to do this: \[ \int x \;dx \]?

OpenStudy (anonymous):

= x^2/2

OpenStudy (anonymous):

Okay. Well \[ \int \frac{2x+1}{3}dx = \frac{1}{3}\left( \int2x\;dx+\int dx\right) \]

OpenStudy (anonymous):

Does it equal to 1/3x^2 ?

OpenStudy (anonymous):

you're missing a term.

OpenStudy (anonymous):

Well what is \[\int\limits dx\]? I'm not so sure because we never learnt it...

OpenStudy (anonymous):

\[ \int dx=x+C \]

OpenStudy (anonymous):

So the answer is \[\frac{ 1 }{ 3 }x^{2}\] ?

OpenStudy (anonymous):

If \(f(x)dx= dx\) then what is \(f(x)\)?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

Oops make that 1/3x^2 + x

OpenStudy (anonymous):

+C

OpenStudy (anonymous):

Ok thanks so much!